OpenStudy (anonymous):

will fan and medal im horrible with math Question 2.2. Write a related equation. k + 26 = 61 (Points : 1) k = 26 61 k = 61 26 k = 61 ÷ 26 k = 61 26

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

C

OpenStudy (anonymous):

i Think

OpenStudy (anonymous):

yyeah i would say the 3rd 1

OpenStudy (anonymous):

thank youu

OpenStudy (anonymous):

Fan And Medal?

OpenStudy (anonymous):

what about Question 3.3. Write a related equation that has the variable on one side. Then simplify the other side. 18b = 18 (Points : 1) b = 0 b = 36 b = 324 b = 1 Question 4.4. Write a related equation that has the variable on one side. Then simplify the other side. (Points : 1) m = 64 m = 4 m = 12 m = 20 Question 5.5. Write a related equation that has the variable on one side. Then simplify the other side. j 5 = 61 (Points : 1) j = 56 j = 12.2 j = 66 j = 305 and yes ill fan & medal thunder

OpenStudy (anonymous):

1 b = 0

OpenStudy (anonymous):

for ??????? 4.4. do you have more backround on it

OpenStudy (anonymous):

61/5=12.2

OpenStudy (anonymous):

5=12.2

OpenStudy (anonymous):

thats all they gave me

OpenStudy (anonymous):

j=49 but not an answer

OpenStudy (anonymous):

alright thanks a bunch

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

if i had an answer it would be j = 12.2

OpenStudy (anonymous):

alright thank you

OpenStudy (lyrae):

k + 26 = 61 k + 26 - 26 = 61 - 26 k = 61 - 26 \[18b = 18\]\[\frac{ 18b }{ 18 } = \frac{ 18 }{ 18 }\]\[b = 1\] 4.4 is missing information. Cannot se the operator between j and 5 in the last one.